### v1

@@ Abstract
Gravitational microlensing occurs when the Schwarzschild radius of the lensing object is smaller or nearly equal to the wavelength of the incoming waves and can produce a chromatic amplitude and phase modulation in the Gravitational Waves (GWs) signal. In contrast, strong gravitational lensing magnifies the incoming gravitational wave in an achromatic way, but can also lead to a constant phase shift. In reality, an observed GW signal can be lensed by both microlensing and strong lensing. To detect and characterize both microlensed and strong lensed GW events together, we have developed a novel method $\mu$-GLANCE (Micro-Gravitational Lensing Authenticator using Non-modelled Cross-correlation Exploration). In this technique, we calculate the cross-correlation of the residual between different detectors with respect to the best fit and use a generalized template to search for both GW source and lensing parameters. We assign a false alarm rate to each candidate from a statistical viewpoint, depending on how many times the noise cross-correlation matches the strength of the residual cross-correlation of the candidate. We show that for an event with a matched-filtering signal-to-noise ratio (SNR) close to thirty, a residual due to microlensing with an amplitude about $10\%$ of strong lensing magnification $\mu\approx 3.2$ will start to show deviation from noise distribution at more than $68\%$ C.I with the LIGO-Virgo-KAGRA sensitivity for the fourth observation run. If the strong lensing or microlensing amplitude is higher (or lower), a lower (or higher) matched-filtering SNR will be required to identify any deviation. This method provides the first technique to detect both strong lensing and micro-lensing without assuming any specific lensing model, and its application on the current and future GW data can identify events with both chromatic and achromatic lensed scenarios.
@@ 1 Introduction
Gravitational waves are propagating fluctuations in the spacetime metric caused by the time-varying quadrupole moment of the mass distribution of a system. With the help of Laser Interferometer Gravitational-wave Observatory (LIGO), we have confirmed the existence of gravitational waves on September 14, 2015 [1]. Until the third observational run (O3) in 2021, the LIGO-Virgo-KAGRA detector network (LVK) has more than 90 confirmed detections of GWs [2, 3, 4, 5, 6, 7, 8, 9]. These observations are caused by some of the most catastrophic processes in the universe, such as mergers of compact object binaries. These detections have opened up a plethora of explorable physics regimes, ranging from cosmology (e.g. [10]), compact binary population inferences (e.g. [11]) to probing fundamental physics (e.g. [12]).
@@
Gravitational lensing is the phenomenon where the presence of a massive object dictates the light-like trajectories passing closeby [13, 14, 15, 16, 17]. The massive object bends the path of the incoming gravitational waves, deflecting them from their original trajectories. This causes the waves to interfere, producing amplitude and phase modulation on the outgoing GW waves. Quite analogous to the different regimes of optics with electromagnetic waves, gravitational lensing of GWs can be treated in two different scenarios: (i) Geometric optics: if the Schwarzschild radius ($R^{s}_{lens}=\frac{2GM}{c^{2}}$) of a massive object (called lens hereafter) is much larger than the wavelength of the emitted gravitational waves ($\lambda_{GW}$), i.e. $R^{s}_{lens}\gg\lambda_{GW}$, we can consider geometric optics approximation. We observe a system of multiple images with GWs, with frequency-independent magnification and phase-shifts. This is referred to as strong lensing scenario. (ii) Wave optics: if the Schwarzschild radius of the lens is of comparable size or smaller than the wavelength of the GW $R^{s}_{lens}\leq\lambda_{GW}$, interfering signals produce GWs with frequency-dependent amplitude modulation and phase modulation factors. This is referred to as microlensing scenario.
@@
Gravitational lensing provides a unique way to infer the density profile, shape and dimension of the massive lensing objects. While electromagnetic waves lensing happens in the geometric optics regime (due to their short wavelengths compared to the large astrophysical objects in the universe), chances of observing microlensing effects on them are negligibly small. On the other hand, GWs having larger wavelengths (minimum $\approx 10^{2}$ km to maximum $\approx 10^{14}$ km [18]), microlensing of GWs is feasible. Therefore, inference of the shape, size and mass distribution of dark matter structures can be performed by studying the microlensing signatures on GWs. Using lensing of GWs in the observable frequency spectra across all current and future GW detectors, we can probe properties of massive bodies of $10^{4}M_{\odot}$ to $10^{14}M_{\odot}$ mass [19, 20]. Lensing of GWs offers unique probe constraints on the astrophysical population of high redshift compact objects [21, 22], cosmological parameters [23, 24, 25] to study the dark matter distribution [26, 27], testing General Theory of Relativity [28, 29, 30] and to detect intermediate and primordial black holes [31, 32].
@@
With the current LIGO-Virgo-KAGRA(LVK) sensitivities [33, 5, 34], observation of a lensed GW is very rare, having the probability of occurrence of a few parts in a thousand (around 0.1 - 0.6%) [35, 22, 36]. The lensing search by the LVK collaboration on the O3 data led to no significant evidence of any lensing of GWs [37] using existing methods [38, 39, 40, 41, 42, 43]. A follow-up study on the O3 events also did not found any significant evidence of the observation of a lensing event so far [44]. With improved detector sensitivities in the upcoming O5 run and with around $\approx 1000$ GW events, we expect to witness at least one lensed GW.
@@
In the cold dark matter model, we expect large dark matter halos to have smaller subhaloes inside it [45]. So while the halo overall as a lensing object produces strong lensing effects on the GWs, the subhaloes (or any present substructures) can contribute to microlensing modifications [46, 47, 48, 49]. Such scenarios not only allow us to study the properties of the dark matter (DM) halo itself but also let us understand the substructures present within the DM halo. However, the microlensing modifications can be an order of magnitude smaller than the strong lensing [50], therefore making it very hard to be observed. However, with advancements in the GW detectors and with the inclusion of next-generation GW observatories, we expect to be able to distinguish such minuscule modifications on the waveform. Such lensing modifications can impose false bias on estimation of the GW source intrinsic parameters [51].
@@
In this perspective, we have developed $\mu$-GLANCE, Micro-Gravitational Lensing Authenticator using Non-modelled Cross-correlation Exploration, a technique to detect microlensed gravitational wave signals. The technique essentially relies on the residual cross-correlation among different detectors. Performing cross-correlation on the residuals, helps us understand the common unmodelled features present in the signals as observed in all detectors. To summarize, we perform two distinct parameter estimations on the data: one with the microlensing hypothesis and the other with no microlensing hypothesis. Therefore, in the strong presence of the microlensing in GW signal (reflected by the support of the microlensing parameters in the probability distributions ), we can detect them straightforward. However, if there is no strong support for the microlensing parameters, we shift our focus on the source parameters obtained using the no microlensing hypothesis and construct the best-fit signal. We cross-correlate between the residuals at different detectors and assign a false alarm rate (FAR) associated. If the FAR is less than the threshold FAR, set at $10^{-3}$ per yr, we call the event as an microlensing candidate. On the other side, if there are multiple images of the same GW i.e. in the strong lensing scenario, we apply GLANCE [52] to detect lensed GW signals. We perform cross-correlation on the reconstructed polarizations and its deviation from the noise cross-correlation helps us determine the significance of the event pair being actually lensed.
@@
Recently, there has emerged several search pipelines to detect microlensed GWs. Microlensing, if present in the data, would have some changed characteristics in the spectrogram of the GW. Therefore, deep learning based methods to detect microlensing signatures from the spectrogram has been performed [53]. Microlensing induces a bias to the source parameter estimation. Therefore, to detect microlensing, a joint parameter estimation based technique has also been implemented to obtain the source and microlens parameters together [39]. Also in the presence of a macrolens and microlens substructures inside, strong and microlensing characteristics can be both present on a GW signal. This would enhance the detection of microlensing due to multiple images of the microlensed GW signal; extracting parameters from those multiple signals would allow to better constrain the microlens parameters [54].
@@
In this work, we have presented $\mu$-GLANCE a novel technique to detect chromatically and achromatically lensed gravitational waves.This work is organised as follows: in section 2, we discuss gravitational lensing of GWs and its strong lensing and microlensing regime. Here, we also motivate about the particular lensing amplification model used in this work. We discuss the salient features of $\mu$-GLANCE, its strengths and weaknesses in detecting lensed GWs in section 3. In section 4 we discuss the methodology used and the mathematical framework behind the cross-correlation based technique. We apply the technique on simulated GW data in section 5 and show its performance of detecting microlensed signals. In section 6, we present the effect of variation of the source and lens parameters and how that supports us for a confident microlensing detection. To remove the inference biases on the source parameters caused by microlensing, we jointly estimate the source and lens parameters in section 7. The false alarm rate associated with a microlensing detection is calculated in section 8. In section 9, we discuss the possible caveats and the future prospects of $\mu$-GLANCE in the detection of a microlensed GW.
@@ 2 Salient aspects of $\mu$-GLANCE
Figure 1: The figure shows the outline for microlensing searches of GWs using $\mu$-GLANCE. We perform two parameter estimation (PE) on the data, one with lensing hypothesis and one without lensing hypothesis. If we observe lensing parameters showing strong support, we can confirm that there is a microlensing signature present in the signal. If not, we perform cross-correlation on the residuals. The residuals are constructed by subtracting the best-fit strain from the data. The best-fit is obtained by using the medians from the distributions from the PE using non-lensing hypothesis.
@@
To look for both strong lensing and microlensing signatures present in GW data, we have developed $\mu$-GLANCE. This method makes both detection and characterization of the lensing signal from the GW data. We show a flowchart describing various parts of $\mu$-GLANCE in figure 1 and describe some the salient aspects of the method below:
  • $\mu$-GLANCE follows a model independent microlensing search approach towards the data. Using Bayesian parameter estimation technique under a non-lensing hypothesis, the source parameters are constrained to obtain a best-fit strain. The best-fit strain is used to obtain the residual from the data. Cross-correlation is performed on the residuals to observe any non-zero deviation, which can be inferred as a microlensing signature present on the data.
@@
  • Lensing modulations can apply bias on the source parameter estimation. Therefore, to remove lensing biases, $\mu$-GLANCE performs a joint source properties and lens characteristics estimation. We construct a model to capture the microlensing amplification effects caused by an astrophysical object on the GW. Therefore, in contrary to strong lensing searches as in GLANCE [52], the microlensing search is not completely unmodelled. We emphasize that the microlensing parameters show up in the joint distribution when there are microlensing modulation present at significant strength. If the modulations are weakly present, the joint source and lens parameter distribution may not be able to constraint the microlensing parameters correctly and in such cases, we depend on the residual cross-correlation signal to detect any microlensing-like signatures.
Also, we need accurate waveform models to capture all features of the GW. For our current analysis use IMRPhenomXPHM [55] waveform model that incorporates eccentric, precessing black hole binary and calculates the waveform for higher modes starting from $(l,m)=(2,2)$  [^1] We have considered the [(2, 2), (2, 1), (3, 3), (3, 2), (4, 4)] and the corresponding negative $m$-modes for the generation of the GW-signal. The inclusion of higher order modes, though marginal in GW data analysis, would have significant impact in the next detectors such as Cosmic Explorer [56] and Einstein Telescope [57]. . In future more accurate waveform model will be required to distinguish between effects arising from waveform mismatch and micro-lensing effects.
3 Basics of Gravitational Lensing: in Microlensing and Strong Lensing Regime

### v2

@@ Abstract
Gravitational microlensing in the wave-optics (WO) regime occurs when the Schwarzschild radius of a lensing object is comparable to or smaller than the wavelength of incoming gravitational waves (GWs), producing chromatic amplitude and phase modulations. In contrary, geometric optics effects happen when wavelength is much smaller than the lensing object leading to frequency independent amplifications and phase shifts. GWs can undergo both effects of lensing due to interactions with objects of different scales. To detect and characterize the wave-optics features from a lensed GW, we have developed a novel method $\mu$-GLANCE (Micro-Gravitational Lensing Authenticator using Non-modelled Cross-correlation Exploration). In this technique, we calculate the cross-correlation between the residuals from different detectors. We assign a false alarm rate to each potential WO microlensed candidate from a statistical viewpoint, depending on how many times the noise cross-correlation matches the strength of the residual cross-correlation of the candidate over a certain period of time. We show that for an event with a matched-filtering signal-to-noise ratio (SNR) close to thirty, a residual due to wave-optics lensing with an amplitude about $10\%$ of magnification $\mu\approx 3.2$ will start to show deviation from noise distribution at more than $68\%$ Confidence interval with the LIGO-Virgo-KAGRA sensitivity for the fourth observation run. This method provides the first technique to detect geometric optics and wave-optics effects from a WO microlensed GW without assuming any specific lensing model, and its application on the current and future GW data can identify events with both chromatic and achromatic lensed scenarios.
@@ 1 Introduction
Gravitational waves (GWs) are propagating fluctuations in the spacetime metric caused by the time-varying quadrupole moment of the mass distribution of a system. With the help of Laser Interferometer Gravitational-wave Observatories (LIGOs) present at Livingston and Hanford sites, we have confirmed the existence of GWs on September 14, 2015 (Abbott et al., 2016a). Until the third observational run (O3) in 2021, the LIGO-Virgo-KAGRA detector network (LVK) has more than 90 confirmed detections of GWs (Abbott et al., 2019, 2021a, 2023a, 2016b; Acernese et al., 2015; Aasi et al., 2015; Buikema et al., 2020a; Akutsu et al., 2021). These observations are caused by some of the most catastrophic processes in the universe, such as mergers of compact object binaries. These detections have opened up a plethora of explorable physics regimes, ranging from cosmology (e.g. (Abbott et al., 2023b)), compact binary population inferences (e.g. (Abbott et al., 2023c)) to probing fundamental physics (e.g. (Abbott et al., 2021b)).
@@
Gravitational lensing is the phenomenon where the presence of a massive object dictates the light-like trajectories passing close-by (Schneider et al., 1992; Wang et al., 1996; Nakamura, 1998; Bartelmann, 2010; Dai et al., 2017). The massive object bends the path of the incoming GWs, deflecting them from their original trajectories. This causes the waves to interfere, producing amplitude and phase modulation on the outgoing GW waves. Quite analogous to the different regimes of optics with electromagnetic waves, gravitational lensing of GWs can be treated in two different scenarios: (i) Geometric optics : if the Schwarzschild radius ($R^{s}_{lens}=\frac{2GM}{c^{2}}$) of a massive object (called lens hereafter) is much larger than the wavelength of the emitted GWs ($\lambda_{GW}$), i.e. $R^{s}_{lens}\gg\lambda_{GW}$, we can consider geometric-optics approximation. We observe a system of multiple images with GWs, with frequency-independent magnification and phase-shifts. (ii) Wave optics (WO) : if the Schwarzschild radius of the lens is of comparable size or smaller than the wavelength of the GW $R^{s}_{lens}\leq\lambda_{GW}$, interfering signals produce GWs with frequency-dependent amplitude modulation and phase modulation factors.
@@
Gravitational lensing provides a unique way to infer the density profile, shape and dimension of the massive lensing objects. While electromagnetic waves lensing happens in the geometric-optics regime (due to their short wavelengths compared to the large astrophysical objects in the universe), chances of observing wave optics effects on them are negligibly small. On the other hand, GWs having larger wavelengths (minimum $\approx 10^{2}$ km to maximum $\approx 10^{14}$ km (Bailes et al., 2021)), WO microlensing of GWs is feasible. Therefore, inference of the shape, size and mass distribution of dark matter structures can be performed by studying the wave-optics lensing signatures on GWs. Using lensing of GWs in the observable frequency spectra across all current and future GW detectors, we can probe properties of massive bodies of $10^{4}M_{\odot}$ to $10^{14}M_{\odot}$ mass (Wambsganss, 2006; Bayer et al., 2023). Lensing of GWs offers unique probe constraints on the astrophysical population of high redshift compact objects Mukherjee et al. (2021a, b), cosmological parameters (Sereno et al., 2011; Balaudo et al., 2023; Jana et al., 2023) to study the dark matter distribution (Massey et al., 2010; Basak et al., 2022), testing General Theory of Relativity (Fan et al., 2017; Mukherjee et al., 2020a, b) and to detect intermediate and primordial black holes (Lai et al., 2018; Oguri & Takahashi, 2020).
@@
With the current LIGO-Virgo-KAGRA(LVK) sensitivities (Abbott et al., 2016c, b; Buikema et al., 2020b), observation of a lensed GW is very rare, having the probability of occurrence of a few parts in a thousand (around 0.1 - 0.6%) (Li et al., 2018; Mukherjee et al., 2021b; Diego et al., 2021). The lensing search by the LVK collaboration on the O3 data led to no significant evidence of any lensing of GWs (The LIGO Scientific Collaboration et al., 2023) using existing methods (Janquart et al., 2021; Wright & Hendry, 2021; Lo & Magaña Hernandez, 2023; Hannuksela et al., 2019; Dai et al., 2020; Singer et al., 2019). A follow-up study on the O3 events also did not found any significant evidence of the observation of a lensing event so far (Janquart et al., 2023). With improved detector sensitivities in the upcoming O5 run and with around $\approx 1000$ GW events, we expect to witness at least one lensed GW.
@@
In the cold dark matter model, we expect large dark matter halos to have smaller sub-haloes inside it (Zavala & Frenk, 2019). So while the halo overall as a lensing object produces geometric-optics effects on the GWs, the sub-haloes (or any present substructures) can contribute to WO microlensing modifications (Fairbairn et al., 2023; Diego et al., 2019; Diego, 2020; Meena, 2024). Such scenarios not only allow us to study the properties of the dark matter (DM) halo itself but also let us understand the substructures present within the DM halo. However, the WO microlensing modifications can be an order of magnitude smaller than the geometric-optics regime (Takahashi & Nakamura, 2003), therefore making it very hard to be observed. However, with advancements in the GW detectors and with the inclusion of next-generation GW observatories, we expect to be able to distinguish such minuscule modifications on the waveform. Such lensing modifications can impose false bias on estimation of the GW source intrinsic parameters (Mishra et al., 2024).
@@
In this perspective, we have developed $\mu$-GLANCE, Micro-Gravitational Lensing Authenticator using Non-modelled Cross-correlation Exploration, a technique to detect WO microlensed GW signals. The technique essentially relies on the residual cross-correlation among different detectors. Performing cross-correlation on the residuals, helps us understand the common unmodelled features present in the signals as observed in all detectors. To summarize, we perform two different Bayesian parameter estimations on the data: one with the WO microlensing hypothesis and the other with no lensing hypothesis. Therefore, in the strong presence of the WO microlensing in GW signal (reflected by the support of the WO microlensing parameters in the probability distributions), we can detect them straightforwardly. However, if there is no strong support for the WO microlensing parameters, we shift our focus on the source parameters obtained using the no WO microlensing hypothesis and construct the best-fit signal. We cross-correlate between the residuals at different detectors and assign a false alarm rate (FAR) associated. If the FAR is less than the threshold FAR, set at $10^{-3}$ per yr, we call the event as an WO microlensed candidate. On the other side, if there are multiple images of the same GW i.e. in the geometric-optics scenario, we apply GLANCE (Chakraborty & Mukherjee, 2024) to detect lensed GW signals. We perform cross-correlation on the reconstructed polarizations and its deviation from the noise cross-correlation helps us determine the significance of the event pair being actually lensed.
@@
Recently, there has emerged several search pipelines to detect WO microlensed GWs. WO microlensing, if present in the data, would have some changed characteristics in the spectrogram of the GW. Therefore, deep learning based methods to detect WO microlensing signatures from the spectrogram has been performed (Kim et al., 2022). WO microlensing induces a bias to the source parameter estimation. Therefore, to detect WO microlensing, a joint parameter estimation based technique has also been implemented to obtain the source and lensing template parameters together (Wright & Hendry, 2021). Also in the presence of a macrolens and microlens substructures inside, geometric-optics and WO effects can be both present on a GW signal. This would enhance the detection of WO microlensing due to multiple images of the WO microlensed GW signal; extracting parameters from those multiple signals would allow to better constrain the WO microlensing parameters (Seo et al., 2022).
@@
In this work, we have presented $\mu$-GLANCE a novel technique to detect chromatically and achromatically lensed GWs. This work is organised as follows: in section 2, we discuss gravitational lensing of GWs and its geometric-optics and WO microlensing regime. Here, we also motivate about the particular lensing amplification model used in this work. We discuss the salient features of $\mu$-GLANCE, its strengths and weaknesses in detecting lensed GWs in section 3. In section 4 we discuss the methodology used and the mathematical framework behind the cross-correlation based technique. We apply the technique on simulated GW data in section 5 and show its performance of detecting WO microlensed signals. In section 6, we present the effect of variation of the source and lens parameters and how that supports us for a confident WO microlensing detection. To remove the inference biases on the source parameters caused by WO microlensing, we jointly estimate the source and lensing template parameters in section 7. The false alarm rate associated with a WO microlensing detection is calculated in section 8. In section 9, we discuss the possible caveats and the future prospects of $\mu$-GLANCE in the detection of a WO microlensed GW.
@@ 2 Salient aspects of $\mu$-GLANCE
$\mu$-GLANCE can look for vast scenarios of WO microlensing signatures present in the data through residual cross-correlation. It can both detect and characterize the lensing signal from the GW data. We show a flowchart describing various parts of $\mu$-GLANCE in figure 1 and describe some of the salient aspects of the method below:
  1. 1.
    Model-independent nature of the algorithm: Any detection method requires to be model-independent or support a large variety of models. Given the limited class of only spherically symmetric lenses studied so far, we cannot make a model-dependent claim of WO microlensing detection. That is why $\mu$-GLANCE follows a first-in-class model-independent approach for WO microlensing search from the data. Through cross-correlation between the residuals from different detectors, $\mu$-GLANCE performs cross-correlation to look for the common non-localized WO microlensing features buried in the data.
  2. 2.
    Removal of noise artifacts though cross-correlation: The detection method needs to be very robust, and it should suppress noise artifacts. Cross-correlation-based $\mu$-GLANCE searches for the presence of any underlying similarities/patterns even at low strengths, while uncorrelated noise from the different detectors is suppressed heavily. Here, we compare the strength of the event cross-correlation as compared to the background noise fluctuations. This helps us understand the significance of an event. Any non-Gaussianity present in the data e.g., glitches which are not correlated among the two detectors, thus $\mu$-GLANCE can handle these situations well.
  3. 3.
    Allowance for different detector pairs for a stronger claim on detection:The detection should have the strength of incorporating multiple detectors for a stronger detection claim. Cross-correlation picks up any similar feature present between the residual pair from a pair of detectors. $\mu$-GLANCE combines all combinations of the pair of detectors and makes a stronger claim for detection if a common feature is observed in different pairs of detectors.
  4. 4.
    Minimization of waveform systematics: Residual morphology depends on the chosen waveform model used for the best-fit subtraction from data. As a result, waveform-based systematics can pose a potential threat to WO microlensing detection. This is why we have made $\mu$-GLANCE in a way that can minimize waveform-based artifacts in the residual cross-correlation. Thus, to check for waveform systematics, $\mu$-GLANCE can perform cross-correlation on residuals subtracting best-fit templates from different waveform models. It helps to benchmark our confidence with the results coming from a specific model and how much it can vary by the changing waveform model. On top of that, by cautious selection of only the part of the waveform from an fmin to fmax in the range over which the waveform systematics are limited. This is built in $\mu$-GLANCE so that it only considers this well-modelled part of the waveform, which is least affected by waveform systematics  [^1] We would like to mention the necessity of very accurate waveform models to capture all features of the GW. For our current analysis, use IMRPhenomXPHM (Pratten et al., 2021) waveform model that incorporates eccentric, precessing black hole binary and calculates the waveform for higher modes starting from $(l,m)=(2,2)$. We have considered the [(2, 2), (2, 1), (3, 3), (3, 2), (4, 4)] and the corresponding negative $m$-modes for the generation of the GW-signal. The inclusion of higher order modes, though marginal in GW data analysis with LVK, would have a significant impact in the next detectors such as Cosmic Explorer (Evans et al., 2021) and Einstein Telescope (Maggiore et al., 2020). .
  5. 5.
@@
    Joint inference of source properties and lensing template properties - to remove WO microlensing biases: WO microlensing modulations are frequency-dependent, often mimicking the effects of GW source properties such as spin, eccentricity, etc. Therefore, it can apply bias to the inference of the source properties. Therefore, to remove lensing biases, $\mu$-GLANCE performs a joint exploration of source properties and lensing template parameters. We construct a very generic model that captures the WO microlensing amplification effects caused by an astrophysical object on the GW. Therefore, in comparison to geometric-optics lensing searches as in GLANCE (Chakraborty & Mukherjee, 2024), the WO microlensing search is not completely unmodelled. The WO microlensing parameters show up in the joint source and lensing template parameters’ distributions when there are WO microlensing modulations present at significant strength. If the modulations are weakly present, the joint source and lensing template parameters distribution may not be able to constrain the WO microlensing parameters well, and in such cases, we solely rely on the residual cross-correlation signal to detect any WO microlensing-like signatures.
Figure 1: The figure shows the outline for WO microlensing searches of GWs using $\mu$-GLANCE. For targeted WO microlensing searches on GW events (The targetted search is performed on a GW event, whose the sky localization is of the order of a few sq. degrees upto tens of sq. degrees consisting of lensed galaxies of relevant redshifts), we first perform the joint source and lensing template parameter estimation (PE) on the data, for the events with high matched-filtering SNR. If we observe strong support for the lensing parameters, we can directly claim the event to be a potential WO microlensing candidate. If no strong support is shown towards the lensing parameters, it is tested though residual cross-correlation in $\mu-$GLANCE. The technique also incorporates any low matched-filtering SNR events to perform a blind WO microlensing search from the data. It performs cross-correlation on the residuals from different detectors. The residuals are constructed by subtracting the best-fit strain (in the unlensed hypothesis) from the data. After the cross-correlation strength has been compared against the noise cross-correlation though its false alarm rate (FAR), a significance of the event for its WO microlensing candidature is evaluated. If the event qualifies strong support though the residual cross-correlation, it is declared as WO microlensing candidate. If the joint parameter estimation has not been performed already, we perform it to constrain the lensing effects on the waveform along with the characterization of the source. A confident detection is approved only when the lensing parameters show deviation from the null lensing hypothesis regime.
These salient aspects of this method make it useful to detect lensing signals from different lensing scenarios, along with mitigating noise contaminations. By this technique, the presence of both WO and geometric-optics lensing signature can be measured from the data, along with distinguishing from unlensed events with a false alarm rate. A detailed discussion on this is presented in Sec. 8. For the scenarios of only geometric-optics lensing and without WO effects (multiple-image system), the performance of this technique is shown by Chakraborty & Mukherjee (2024).
2.1 Comparison with the existing WO microlensing-based searches:
This new method $\mu$-GLANCE estimates the best-fit unlensed waveform to obtain the residuals from each detector for a GW event. These residuals are then studied through cross-correlation for any common signature not captured in the waveform model. The relative strength of the cross-correlation signal as compared to the noise fluctuations helps us quantify the detection significance. This $\mu$-GLANCE is the very first model-independent search technique to look for WO microlensing signatures. From the entire third run of the LVK observatories, the LVK collaboration searched for potential WO microlensed signals from the data (Abbott et al., 2021c, 2024). It estimated the Bayes factor, which shows the support for the lensing hypothesis as compared to the non-lensing hypothesis by comparing the respective likelihoods. The work observes no strong evidence for any WO microlensed candidate by observing the Bayes factors. However, the work has considered point-mass lensing objects for their analysis. The point-mass lens is not a good approximation for its spherical symmetry, a very uncommon trait for an astronomical object.
Recently, machine-learning-based techniques (Kim et al., 2022) have been applied to the data to search for point-mass lens effects on the data. These methods use spectrogram (q-transformed data, which maps the evolution of the GW event through the time-frequency bins). However, such searches are not conclusive since the mass distribution of lensing objects is most of the time not spherically symmetric. Also, the waveform model for the training dataset is IMRPhenomPv2, which is not inclusive of all such different models.
Joint estimations of the source and the lensing template parameters are also performed for WO microlensing detection in a previous work (Wright & Hendry, 2021). The work incorporates the wave-optics effects of the GW for point-mass, singular-isothermal-sphere (SIS), and Navarro-Frenk-White (NFW) mass distribution profiles. However, these distributions are all spherically symmetric, for objects of sizes of the GW wavelength, the assumption of the spherical symmetry of the lens cannot be confirmed, Therefore, a more rigorous way would either require sophisticated lens profiles or need a completely model-free approach, which is not followed in this technique.
3 Basics of Gravitational Lensing: in Wave-optics and Geometric-optics lensing Regimes